# Menger and consonant sets in the Sacks model

### v2

@@ 1. Introduction
By space we mean an infinite topological Tychonoff space. A space $X$ is Menger if for any sequence $\mathcal{U}_{0},\allowbreak\mathcal{U}_{1},\allowbreak\dotsc$ of open covers of $X$, there are finite families $\mathcal{F}_{0}\subseteq\mathcal{U}_{0},\mathcal{F}_{1}\subseteq\mathcal{U}_{1},\dotsc$ such that the family $\bigcup_{n\in\omega}\mathcal{F}_{n}$ covers $X$. Every $\sigma$-compact space is Menger and every Menger space is Lindelöf. The Menger conjecture asserts that every subset of the real line with the Menger property is $\sigma$-compact. By a result of Fremlin and Miller [13, Theorem 4], this conjecture is false in ZFC. This opened a wide stream of investigations in the realm of special subsets of the real line. The Menger property is a subject of research in the combinatorial covering properties theory but also appears in other branches of mathematics as local properties of function spaces [16], forcing theory [10, 11] or additive Ramsey theory in algebra [30, 26].
The Menger property is closely related to infinite combinatorics. Let $a,b\in{\omega^{\omega}}$. We write $a\leq^{*}b$ if the set $\{\,n:a(n)>b(n)\,\}$ is finite. In such a case we say that the function $a$ is dominated by the function $b$. A set $D\subseteq{\omega^{\omega}}$ is dominating if any function in ${\omega^{\omega}}$ is dominated by some function from $D$. Let $\mathfrak{d}$ be the minimal cardinality of a dominating subset of ${\omega^{\omega}}$. The Menger property can be characterized in terms of continuous images, as follows: a set $X\subseteq{2^{\omega}}$ is Menger if and only if no continuous image of $X$ into ${\omega^{\omega}}$ is dominating. This characterization was proved by Hurewicz [14, § 5] and then much later but independently by Recław [21, Proposition 3], so we call it the Hurewicz-Recław characterization of the Menger property. It follows that any subset of $2^{\omega}$ with cardinality smaller than $\mathfrak{d}$ is Menger and there is a non-Menger set of cardinality $\mathfrak{d}$.
@@
Theorem 1.1 (Bartoszyński, Tsaban [3, Theorem 16]).
@@
In the second part of the paper we analyze the structure of Hurewicz and consonant subsets of ${2^{\omega}}$. Recall that a space $X$ is Hurewicz if for any sequence $\mathcal{U}_{0},\allowbreak\mathcal{U}_{1},\allowbreak\dotsc$ of open covers of $X$, there are finite families $\mathcal{F}_{0}\subseteq\mathcal{U}_{0},\mathcal{F}_{1}\subseteq\mathcal{U}_{1},\dotsc$ such that the family $\{\bigcup\mathcal{F}_{n}:n\in\omega\}$ is a $\gamma$-cover of $X$, i.e., the sets $\{\,n:x\in\bigcup\mathcal{F}_{n}\,\}$ are cofinite for all $x\in X$. Obviously, every $\sigma$-compact space is Hurewicz and every Hurewicz space is Menger. Similarly to the Menger property, the Hurewicz property can be characterized in terms of continuous images, as follows. A set $A\subseteq{\omega^{\omega}}$ is bounded if there is a function $b\in{\omega^{\omega}}$ such that $a\leq^{*}b$ for all $a\in A$. A set $X\subseteq{2^{\omega}}$ is Hurewicz if and only if every continuous image of $X$ into ${\omega^{\omega}}$ is bounded. This characterization was proved independently by Hurewicz [14, § 5] and Recław [21, Proposition 1], so we again call it the Hurewicz-Recław characterization of the Hurewicz property.
Consonant spaces were introduced by Dolecki, Greco and Lechicki [12] and characterized by Jordan [15, Corollary 11] in the following way. Let $X\subseteq{2^{\omega}}$. A cover of $X$ is a k-cover if any compact subset of $X$ is contained in some set from the cover. A game $\mathsf{G}_{1}(\mathcal{K},\mathcal{O})$ played on $X$ is a game for two players, Alice and Bob. For a natural number $n$, in round $n$: Alice picks an open $k$-cover $\mathcal{U}_{n}$ of $X$ and Bob picks a set $U_{n}\in\mathcal{U}_{n}$. Bob wins the game if the family $\{\,U_{n}:n\in\omega\,\}$ is a cover of $X$, and Alice wins otherwise. A set $Y\subseteq 2^{\omega}$ is consonant if and only if Alice has no winning strategy in the game $\mathsf{G}_{1}(\mathcal{K},\mathcal{O})$ played on $2^{\omega}\setminus Y$. We treat here this characterization as a definition of consonant sets.
@@
In Section 5 we introduce a topological game which allows us to analyze the structure of Hurewicz and consonant spaces in ${2^{\omega}}$. We show that in the Sacks model, mentioned above, each consonant (Hurewicz) set $X\subseteq{2^{\omega}}$ and its complement ${2^{\omega}}\setminus X$ are unions of $\omega_{1}$ compact sets. This approach allows us also to capture a result of Miller [17, §5] that in this model, every perfectly meager subset of ${2^{\omega}}$, i.e., a set which is meager in any perfect subset of ${2^{\omega}}$, has size at most $\omega_{1}$.
The main tools used in our investigations are game characterizations of the considered properties. Let $X$ be a space. The Menger game played on $X$ is a game for two players, Alice and Bob. For a natural number $n$, in round $n$: Alice picks an open cover $\mathcal{U}_{n}$ of $X$ and Bob picks a finite family $\mathcal{F}_{n}\subseteq\mathcal{U}_{n}$. Bob wins the game if the family $\bigcup_{n\in\omega}\mathcal{F}_{n}$ covers $X$, and Alice wins otherwise. For more details about this game, we refer to the works of Scheepers [24, Theorem 13] or Tsaban and the second named author [27]. Similarly to the Menger property, the Hurewicz property can also be characterized using a topological game. The Hurewicz game played on $X$ is a game for two players, Alice and Bob. For a natural number $n$, in round $n$: Alice picks an open cover $\mathcal{U}_{n}$ of $X$ and Bob picks a finite family $\mathcal{F}_{n}\subseteq\mathcal{U}_{n}$. Bob wins the game if the family $\{\,\bigcup\mathcal{F}_{n}:n\in\omega\,\}$ is a $\gamma$-cover of $X$, and Alice wins otherwise.
@@
Theorem 1.5 (Scheepers [24, Theorem 27]).
@@ 2. Menger game and perfect sets
$[\nu]:=\{\,x\in({2^{\omega}})^{S}:x(\beta)\upharpoonright k_{n}=\nu(\beta)\text{ for all }\beta\in F_{n}\,\}.$
@@
Fix $n$ and enumerate $\Sigma_{n}$ injectively as $\{\nu_{0},\ldots,\nu_{N}\}$. Since $[\nu_{0}]\cap K$ is perfect by $(e_{f})$, we can pick $E_{0}(\nu_{0})\in([\nu_{0}]\cap K)\setminus X$. Fix a number $k<N$ and put $C:=\{\nu_{i}:i\leq k\}$. Assume that we have already defined a coherent map $E_{k}:C\to K\setminus X$. Then for every $m>n$ and $\nu_{i}\in C$, there is $e^{m}_{0}(\nu_{i})\in\Sigma_{m}$ such that $E_{k}(\nu_{i})\in[e^{m}_{0}(\nu_{i})]$. In that way we define a map $e^{m}_{0}\colon C\to\Sigma_{m}$. It is clear that this map is coherent. By $(e_{f})$ there are $\mu_{\langle 0\rangle}\neq\mu_{\langle 1\rangle}\in\Sigma_{n+1}$ such that both $e^{n+1}_{\langle 0\rangle}=e^{n+1}_{0}\cup\{\langle\nu_{k+1},\mu_{\langle 0\rangle}\rangle\}$ and $e^{n+1}_{\langle 1\rangle}=e^{n+1}_{0}\cup\{\langle\nu_{k+1},\mu_{\langle 1\rangle}\rangle\}$ are coherent as maps from $C\cup\{\nu_{k+1}\}$ to $\Sigma_{n+1}$.
@@
We shall describe a strategy §$\S$ for Alice in the Menger game played on $X$ such that each play lost by Alice gives rise to the objects whose existence we need to establish. For every $n\in\omega$, let $E_{n}\colon\Sigma_{n}\to K$ be a coherent selection from Lemma 2.1. Put
@@
where $j\geq i_{0}$, is an increasing open cover of $X$. Now, §$\S$ instructs Alice to start the play with $\mathcal{U}_{0}$.
@@
where $j\geq i_{n}$, is an increasing open cover of $X$. Now, §$\S$ instructs Alice to play the family $\mathcal{U}_{n}$.
@@
This completes our definition of the strategy §$\S$ for Alice such that each infinite play in the Menger game on $X$ in which Alice uses §$\S$ gives rise to a sequence
@@
as described above. In particular, conditions $(1)$$(4)$ are satisfied by the construction. By Theorem 1.4, there is a play, where Alice uses the strategy §$\S$ and the play is won by Bob. Then $X\subseteq\bigcup_{n\in\omega}U^{n}_{j_{n}}$, i.e.,
@@ 3. Combinatorics of conditions in the iterated Sacks forcing
Let $2^{<\omega}:=\bigcup_{n\in\omega}2^{n}$. For elements $s,t\in 2^{<\omega}$, we write $s\subseteq t$ if the sequence $s$ is an initial segment of the sequence $t$, i.e., $s(i)=t(i)$ for all $i\in\operatorname{dom}(s)$. A Sacks tree is a set $p\subseteq 2^{<\omega}$ such that for every $s\in p$ and a natural number $n$, we have $s\restriction n\in p$ and there are elements $t,u\in p$ with $s\subseteq t$, $s\subseteq u$, $t\not\subseteq u$ and $t\not\subseteq u$. For Sacks trees $p$ and $q$, a condition $q$ is stronger than $p$ which we write $q\geq p$ if $q\subseteq p$. The Sacks poset $\mathbb{S}$ is the set of all Sacks trees ordered by $\geq$. For $p,q\in\mathbb{S}$ and natural numbers $m>n$, we write $(q,m)\geq(p,n)$ if $q\subseteq p$ and for every $s\in p\cap 2^{n}$, there are different elements $t,u\in q\cap 2^{m}$ such that $s\subseteq t$ and $s\subseteq u$. For $p\in\mathbb{S}$ and $s\in p$, let $p_{s}:=\{\,t\in p:t\subseteq s\text{ or }s\subseteq t\,\}$.
@@
Lemma 3.3 (Miller [17, Lemma 2]).
@@
Lemma 3.5 (Miller [17, Lemma 5]).
@@
Now we proceed by induction on the cardinality of $F$. If $F=\{\xi\}$, then $|X_{i}|=1$ for all $i<N$ (because $12^{n\cdot(|F|-1)}=$ 1), i.e., $X_{i}=\{x_{i}\}$ for some $x_{i}\in 2^{\omega}$. Put $y_{\mu_{i}}:=x_{i}\upharpoonright l_{*}$ for all $i<N$ and let $r\in\mathbb{S}_{\alpha}$ be a condition such that $r\upharpoonright\xi=u^{l_{*}}_{N-1}\upharpoonright\xi$, $r\upharpoonright\xi$ forces $r(\xi)$ to be $\bigcup\{\,u^{l_{*}}_{i}(\xi):i<N\,\}$, and $(r|\mu_{i})\upharpoonright\beta$ forces $r(\beta)=u^{l_{*}}_{i}(\beta)$ for all $\beta>\xi$. It follows from the above that $r|\mu_{i}\geq u^{l_{*}}_{i}$ and hence $r|\mu_{i}$ forces $\tau\upharpoonright l_{*}=x_{i}\upharpoonright l_{*}=y_{\mu_{i}}$, hence $q:=r$ and $l:=l_{*}$ are as required.
@@
The following fact is reminiscent of [5, Lemma 2.3(i)], and we use a rather similar approach to the proof, which we present for the sake of completeness.
@@
Lemma 3.7 (Miller [17, Lemma 6]).
@@
Let $p$ be the fusion of the sequence ${\langle\,(p_{n},k_{n},F_{n}):n\in\omega\,\rangle}$ [5, Lemma 1.2] and note that it is as required. ∎
@@
We use notation from the formulation of Lemma 3.7. Set $S=\mathrm{supp}(p)$, $\Sigma_{n}=\{\nu\in(2^{k_{n}})^{F_{n}}:\nu$ is consistent with $p\}$, and suppose that ${\langle\,\langle i_{n},j_{n},C_{n}\rangle:n\in\omega\,\rangle}$ is a sequence such that items $(1)$$(4)$ of Lemma 2.2 are satisfied. Then there exists $q\geq p$ such that $\{p|\sigma:\sigma\in C_{n}\}$ is predense above $q$ for all $n\in\omega$.

### v3

@@ 1. Introduction
By space we mean an infinite topological Tychonoff space. A space $X$ is Menger if for any sequence $\mathcal{U}_{0},\allowbreak\mathcal{U}_{1},\allowbreak\dotsc$ of open covers of $X$, there are finite families $\mathcal{F}_{0}\subseteq\mathcal{U}_{0},\mathcal{F}_{1}\subseteq\mathcal{U}_{1},\dotsc$ such that the family $\bigcup_{n\in\omega}\mathcal{F}_{n}$ covers $X$. Every $\sigma$-compact space is Menger and every Menger space is Lindelöf. The Menger conjecture asserts that every subset of the real line with the Menger property is $\sigma$-compact. By a result of Fremlin and Miller [12, Theorem 4], this conjecture is false in ZFC. This opened a wide stream of investigations in the realm of special subsets of the real line. The Menger property is a subject of research in the combinatorial covering properties theory but also appears in other branches of mathematics as local properties of function spaces [15], forcing theory [9, 10] or additive Ramsey theory in algebra [27, 24].
The Menger property is closely related to infinite combinatorics. Let $a,b\in{\omega^{\omega}}$. We write $a\leq^{*}b$ if the set $\{\,n:a(n)>b(n)\,\}$ is finite. In such a case we say that the function $a$ is dominated by the function $b$. A set $D\subseteq{\omega^{\omega}}$ is dominating if any function in $\omega^{\omega}$} is dominated by some function from $D$. Let $\mathfrak{d}$ be the minimal cardinality of a dominating subset of $\omega^{\omega}$}. The Menger property can be characterized in terms of continuous images, as follows: a set $X\subseteq{2^{\omega}}$ is Menger if and only if no continuous image of $X$ into $\omega^{\omega}$} is dominating. This characterization was proved by Hurewicz [13, § 5] and then much later but independently by Recław [19, Proposition 3], so we call it the Hurewicz-Recław characterization of the Menger property. It follows that any subset of $2^{\omega}$ with cardinality smaller than $\mathfrak{d}$ is Menger and there is a non-Menger set of cardinality $\mathfrak{d}$.
@@
Theorem 1.1 (Bartoszyński, Tsaban [2, Theorem 16]).
@@
In the second part of the paper we analyze the structure of Hurewicz and consonant subsets of $2^{\omega}$}. Recall that a space $X$ is Hurewicz if for any sequence $\mathcal{U}_{0},\allowbreak\mathcal{U}_{1},\allowbreak\dotsc$ of open covers of $X$, there are finite families $\mathcal{F}_{0}\subseteq\mathcal{U}_{0},\mathcal{F}_{1}\subseteq\mathcal{U}_{1},\dotsc$ such that the family $\{\bigcup\mathcal{F}_{n}:n\in\omega\}$ is a $\gamma$-cover of $X$, i.e., the sets $\{\,n:x\in\bigcup\mathcal{F}_{n}\,\}$ are cofinite for all $x\in X$. Obviously, every $\sigma$-compact space is Hurewicz and every Hurewicz space is Menger. Similarly to the Menger property, the Hurewicz property can be characterized in terms of continuous images, as follows. A set $A\subseteq{\omega^{\omega}}$ is bounded if there is a function $b\in{\omega^{\omega}}$ such that $a\leq^{*}b$ for all $a\in A$. A set $X\subseteq{2^{\omega}}$ is Hurewicz if and only if every continuous image of $X$ into $\omega^{\omega}$} is bounded. This characterization was proved independently by Hurewicz [13, § 5] and Recław [19, Proposition 1], so we again call it the Hurewicz-Recław characterization of the Hurewicz property.
Consonant spaces were introduced by Dolecki, Greco and Lechicki [11] and characterized by Jordan [14, Corollary 11] in the following way. Let $X\subseteq{2^{\omega}}$. A cover of $X$ is a k-cover if any compact subset of $X$ is contained in some set from the cover. A game $\mathsf{G}_{1}(\mathcal{K},\mathcal{O})$ played on $X$ is a game for two players, Alice and Bob. For a natural number $n$, in round $n$: Alice picks an open $k$-cover $\mathcal{U}_{n}$ of $X$ and Bob picks a set $U_{n}\in\mathcal{U}_{n}$. Bob wins the game if the family $\{\,U_{n}:n\in\omega\,\}$ is a cover of $X$, and Alice wins otherwise. A set $Y\subseteq 2^{\omega}$ is consonant if and only if Alice has no winning strategy in the game $\mathsf{G}_{1}(\mathcal{K},\mathcal{O})$ played on $2^{\omega}\setminus Y$. We treat here this characterization as a definition of consonant sets.
@@
In Section 5 we introduce a topological game which allows us to analyze the structure of Hurewicz and consonant spaces in $2^{\omega}$}. We show that in the Sacks model, mentioned above, each consonant (Hurewicz) set $X\subseteq{2^{\omega}}$ and its complement ${2^{\omega}}\setminus X$ are unions of $\omega_{1}$ compact sets. This approach allows us also to capture a result of Miller [16, §5] that in this model, every perfectly meager subset of $2^{\omega}$}, i.e., a set which is meager in any perfect subset of $2^{\omega}$}, has size at most $\omega_{1}$.
The main tools used in our investigations are game characterizations of the considered properties. Let $X$ be a space. The Menger game played on $X$ is a game for two players, Alice and Bob. For a natural number $n$, in round $n$: Alice picks an open cover $\mathcal{U}_{n}$ of $X$ and Bob picks a finite family $\mathcal{F}_{n}\subseteq\mathcal{U}_{n}$. Bob wins the game if the family $\bigcup_{n\in\omega}\mathcal{F}_{n}$ covers $X$, and Alice wins otherwise. For more details about this game, we refer to the works of Scheepers [22, Theorem 13] or Tsaban and the second named author [25]. Similarly to the Menger property, the Hurewicz property can also be characterized using a topological game. The Hurewicz game played on $X$ is a game for two players, Alice and Bob. For a natural number $n$, in round $n$: Alice picks an open cover $\mathcal{U}_{n}$ of $X$ and Bob picks a finite family $\mathcal{F}_{n}\subseteq\mathcal{U}_{n}$. Bob wins the game if the family $\{\,\bigcup\mathcal{F}_{n}:n\in\omega\,\}$ is a $\gamma$-cover of $X$, and Alice wins otherwise.
@@
Theorem 1.5 (Scheepers [22, Theorem 27]).
@@ 2. Menger game and perfect sets
$\nu]:=\{\,x\in({2^{\omega}})^{S}:x(\beta)\upharpoonright k_{n}=\nu(\beta)\text{ for all }\beta\in F_{n}\,\}.$
@@
Fix $n$ and enumerate $\Sigma_{n}$ injectively as $\{\nu_{0},\ldots,\nu_{N}\}$. Since $\nu_{0}]\cap K$ is perfect by $(e_{f})$, we can pick $E_{0}(\nu_{0})\in([\nu_{0}]\cap K)\setminus X$. Fix a number $k<N$ and put $C:=\{\nu_{i}:i\leq k\}$. Assume that we have already defined a coherent map $E_{k}:C\to K\setminus X$. Then for every $m>n$ and $\nu_{i}\in C$, there is $e^{m}_{0}(\nu_{i})\in\Sigma_{m}$ such that $E_{k}(\nu_{i})\in[e^{m}_{0}(\nu_{i})]$. In that way we define a map $e^{m}_{0}\colon C\to\Sigma_{m}$. It is clear that this map is coherent. By $(e_{f})$ there are $\mu_{\langle 0\rangle}\neq\mu_{\langle 1\rangle}\in\Sigma_{n+1}$ such that both $e^{n+1}_{\langle 0\rangle}=e^{n+1}_{0}\cup\{\langle\nu_{k+1},\mu_{\langle 0\rangle}\rangle\}$ and $e^{n+1}_{\langle 1\rangle}=e^{n+1}_{0}\cup\{\langle\nu_{k+1},\mu_{\langle 1\rangle}\rangle\}$ are coherent as maps from $C\cup\{\nu_{k+1}\}$ to $\Sigma_{n+1}$.
@@
We shall describe a strategy $\S$ for Alice in the Menger game played on $X$ such that each play lost by Alice gives rise to the objects whose existence we need to establish. For every $n\in\omega$, let $E_{n}\colon\Sigma_{n}\to K$ be a coherent selection from Lemma 2.1. Put
@@
where $j\geq i_{0}$, is an increasing open cover of $X$. Now, $\S$ instructs Alice to start the play with $\mathcal{U}_{0}$.
@@
where $j\geq i_{n}$, is an increasing open cover of $X$. Now, $\S$ instructs Alice to play the family $\mathcal{U}_{n}$.
@@
This completes our definition of the strategy $\S$ for Alice such that each infinite play in the Menger game on $X$ in which Alice uses $\S$ gives rise to a sequence
@@
as described above. In particular, conditions (1)(1)$(4)$ are satisfied by the construction. By Theorem 1.4, there is a play, where Alice uses the strategy $\S$ and the play is won by Bob. Then $X\subseteq\bigcup_{n\in\omega}U^{n}_{j_{n}}$, i.e.,
@@ 3. Combinatorics of conditions in the iterated Sacks forcing
[…]
@@
Lemma 3.3 (Miller [16, Lemma 2]).
@@
Lemma 3.5 (Miller [16, Lemma 5]).
@@
Now we proceed by induction on the cardinality of $F$. If $F=\{\xi\}$, then $|X_{i}|=1$ for all $i<N$ (because $2^{n\cdot(|F|-1)}=1$), i.e., $X_{i}=\{x_{i}\}$ for some $x_{i}\in 2^{\omega}$. Put $y_{\mu_{i}}:=x_{i}\upharpoonright l_{*}$ for all $i<N$ and let $r\in\mathbb{S}_{\alpha}$ be a condition such that $r\upharpoonright\xi=u^{l_{*}}_{N-1}\upharpoonright\xi$, $r\upharpoonright\xi$ forces $r(\xi)$ to be $\bigcup\{\,u^{l_{*}}_{i}(\xi):i<N\,\}$, and $(r|\mu_{i})\upharpoonright\beta$ forces $r(\beta)=u^{l_{*}}_{i}(\beta)$ for all $\beta>\xi$. It follows from the above that $r|\mu_{i}\geq u^{l_{*}}_{i}$ and hence $r|\mu_{i}$ forces $\tau\upharpoonright l_{*}=x_{i}\upharpoonright l_{*}=y_{\mu_{i}}$, hence $q:=r$ and $l:=l_{*}$ are as required.
@@
The following fact is reminiscent of [4, Lemma 2.3(i)], and we use a rather similar approach to the proof, which we present for the sake of completeness.
@@
Lemma 3.7 (Miller [16, Lemma 6]).
@@
Let $p$ be the fusion of the sequence $\langle\,(p_{n},k_{n},F_{n}):n\in\omega\,\rangle$} [4, Lemma 1.2] and note that it is as required. ∎
@@
We use notation from the formulation of Lemma 3.7. Set $S=\mathrm{supp}(p)$, $\Sigma_{n}=\{\nu\in(2^{k_{n}})^{F_{n}}:\nu$ is consistent with p}p\}, and suppose that $\langle\,\langle i_{n},j_{n},C_{n}\rangle:n\in\omega\,\rangle$} is a sequence such that items (1)(1)$(4)$ of Lemma 2.2 are satisfied. Then there exists $q\geq p$ such that $\{p|\sigma:\sigma\in C_{n}\}$ is predense above $q$ for all $n\in\omega$.